OpenStudy (anonymous):

Tom wants to hang a mirror in his room. The mirror and frame must have an area of 14 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (jack1):

A = L x W (mirror and frame) do we assume uniform thickness of frame?

OpenStudy (anonymous):

Yes.

OpenStudy (jack1):

\[14 = (L) \times (W)\] \[14 = (Mirror L + Frame L) \times (Mirror W + Frame W)\] \[14 = (4 + x) \times (3 + x)~~\text{if uniform thickness}\] \[\text{use FOIL to expand}~~(4 + x) (3 + x)~~\] can u solve from here @BasedGod1122 ???

OpenStudy (anonymous):

12+4x+3x+x^2 ?

OpenStudy (jack1):

yep, that's ur quadratic, and that = 14

OpenStudy (jack1):

no, that equals 14, if u take 14 from both sides, u can find the zero's...

OpenStudy (anonymous):

x^2 + 7x - 2 = 0 solve for x to find thickness

OpenStudy (anonymous):

How do I solve for x? Walk me through the steps?

OpenStudy (anonymous):

x = 0.27491721763537484861834240347306 feet

OpenStudy (jack1):

do you know the quadratic equation...?

OpenStudy (anonymous):

This is a multiple choice question. My choices are: 4x2 + 14x - 2 = 0 3x2 + 10x - 14 = 0 2x2 + 12x - 14 = 0 x2 + 7x - 14 = 0

OpenStudy (anonymous):

use the formula \[\frac{ -b +- \sqrt{b ^{2}-4ac} } {2}\]

OpenStudy (jack1):

none of the above options work, as @kx2bay said, the quadratic equation comes down to x^2 + 7x - 2 = 0 so non of ur answers match...?

OpenStudy (anonymous):

yes jack is right if you substitute x = 0.27492 in jacks formula above 14=(4+x)×(3+x) u get 14=14

OpenStudy (anonymous):

So which choice do i select

OpenStudy (anonymous):

none of the above choices is correct, I'd say they made a typo with the 4th choice x2 + 7x - 14 = 0 where -14 should be -2 so go with this choice

OpenStudy (anonymous):

Thank you, friend. You've really helped me out. I hope that you have a good night.

OpenStudy (anonymous):

thanks you too, its actually day light here in melbourne australia ;)